# NaturalSpeech 3: Zero-Shot Speech Synthesis with Factorized Codec and Diffusion Models

### v2

@@ Abstract
While recent large-scale text-to-speech (TTS) models have achieved significant progress, they still fall short in speech quality, similarity, and prosody. Considering speech intricately encompasses various attributes (e.g., content, prosody, timbre, and acoustic details) that pose significant challenges for generation, a natural idea is to factorize speech into individual subspaces representing different attributes and generate them individually. Motivated by it, we propose NaturalSpeech 3, a TTS system with novel factorized diffusion models to generate natural speech in a zero-shot way. Specifically, 1) we design a neural codec with factorized vector quantization (FVQ) to disentangle speech waveform into subspaces of content, prosody, timbre, and acoustic details; 2) we propose a factorized diffusion model to generate attributes in each subspace following its corresponding prompt. With this factorization design, NaturalSpeech 3 can effectively and efficiently model the intricate speech with disentangled subspaces in a divide-and-conquer way. Experiments show that NaturalSpeech 3 outperforms the state-of-the-art TTS systems on quality, similarity, prosody, and intelligibility, and achieves on-par quality with human recordings. Furthermore, we achieve better performance by scaling to 1B parameters and 200K hours of training data.
Figure 1: (a) Overview of NaturalSpeech 3, with a neural speech codec for speech attribute factorization and a factorized diffusion model. (b) Data and model scaling of NaturalSpeech 3.
@@ 3.2.1 FACodec Model Overview
We propose a factorized neural speech codec (i.e., FACodec222 We will release the code and pre-trained checkpoint of FACodec soon.) to convert complex speech waveform into disentangled subspaces representing speech attributes of content, prosody, timbre, and acoustic details and reconstruct high-quality speech waveform from these.
@@ 4.4 Effectiveness of Data and Model Scaling
In this section, we study the effectiveness of data and model scaling on the proposed factorized diffusion model. We use the same FACodec trained on LibriLight dataset for fair comparison. We evaluate the zero-shot TTS performance in terms of speaker similarity (Sim-O) and robustness (WER) on an internal test set consisting of $30$ audio clips.
Data Scaling. With a fixed model size of 500M parameters, we trained the factorized diffusion model on three datasets: 1) a 1K-hour subset randomly drawn from the Librilight dataset, 2) a 60K-hour Librilight dataset, and 3) an internal dataset with 200K hours of speech. In Table 7, we observe that: 1) even with a mere 1K hours of speech data, our model attains a Sim-O score of $0.69$ and a WER of $3.39$. It shows that with the speech factorization, NaturalSpeech 3 can generate the speech effectively. 2) As we scale up training data from 1K hours to 60K hours, and then to 200K hours, NaturalSpeech 3 displays continuously enhanced performance, with an improvement of $0.03$ and $0.04$ in terms of Sim-O, and $0.33$ and $0.56$ in terms of WER, respectively, thus confirming the benefits of data scaling. Note that our method trained on 200K hours is still underfitting and longer training will result in better performance.
Model Scaling. We scale up the model size from 500M to 1B parameters with the internal 200K hours dataset. Specifically, we double the number of transformer layers from $12$ to $24$. The results in Table 8 show a boost in both speaker similarity ($0.02$ in Sim-O) and robustness ($0.21$ in WER), validating the effectiveness of model scaling. In the future, we will scale up the model size even larger to achieve better results.

### v3

@@ Abstract
While recent large-scale text-to-speech (TTS) models have achieved significant progress, they still fall short in speech quality, similarity, and prosody. Considering speech intricately encompasses various attributes (e.g., content, prosody, timbre, and acoustic details) that pose significant challenges for generation, a natural idea is to factorize speech into individual subspaces representing different attributes and generate them individually. Motivated by it, we propose NaturalSpeech 3, a TTS system with novel factorized diffusion models to generate natural speech in a zero-shot way. Specifically, 1) we design a neural codec with factorized vector quantization (FVQ) to disentangle speech waveform into subspaces of content, prosody, timbre, and acoustic details; 2) we propose a factorized diffusion model to generate attributes in each subspace following its corresponding prompt. With this factorization design, NaturalSpeech 3 can effectively and efficiently model intricate speech with disentangled subspaces in a divide-and-conquer way. Experiments show that NaturalSpeech 3 outperforms the state-of-the-art TTS systems on quality, similarity, prosody, and intelligibility, and achieves on-par quality with human recordings. Furthermore, we achieve better performance by scaling to 1B parameters and 200K hours of training data.
Refer to captionRefer to caption
Figure 1: (a) Overview of the system, with a neural speech codec for speech attribute factorization and a factorized diffusion model. (b) Data and model scaling of the system.
@@ 3.2.1 FACodec Model Overview
We propose a factorized neural speech codec (i.e., FACodec222 We release the code and pre-trained checkpoint of FACodec at https://huggingface.co/spaces/amphion/naturalspeech3_facodec.) to convert complex speech waveform into disentangled subspaces representing speech attributes of content, prosody, timbre, and acoustic details and reconstruct high-quality speech waveform from these.
@@ 4.4 Effectiveness of Data and Model Scaling
In this section, we study the effectiveness of data and model scaling on the proposed factorized diffusion model. We evaluate the zero-shot TTS performance in terms of speaker similarity (Sim-O) and robustness (WER) on an internal test set consisting of $30$ audio clips.
Data Scaling. With a fixed model size of 500M parameters, we train NaturalSpeech 3, including both FACodec and the factorized diffusion model, on three datasets: 1) a 1K-hour subset randomly drawn from the Librilight dataset, 2) a 60K-hour Librilight dataset, and 3) an internal dataset with 200K hours of speech. In Table 7, we observe that: 1) even with a mere 1K hours of speech data, our model attains a Sim-O score of $0.64$ and a WER of $3.94$. It shows that with the speech factorization, NaturalSpeech 3 can generate the speech effectively. 2) As we scale up training data from 1K hours to 60K hours, and then to 200K hours, NaturalSpeech 3 displays continuously enhanced performance, with an improvement of $0.08$ and $0.09$ in terms of Sim-O, and $0.91$ and $1.83$ in terms of WER, respectively, thus confirming the benefits of data scaling.
Model Scaling. We scale up the model size of the factorized diffusion model from 500M to 1B parameters with the internal 200K hours dataset. Specifically, we double the number of transformer layers from $12$ to $24$. The results in Table 8 show a boost in both speaker similarity ($0.05$ in Sim-O) and robustness ($0.40$ in WER), validating the effectiveness of model scaling. In the future, we will scale up the model size even larger to achieve better results.
5 Conclusion